THE ORIGIN BAND OF THE b̃ − ã SYSTEM OF CH_{2}

GREGORY E. HALL, AND TREVOR J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, NY, 11973; BOR-CHEN CHANG, Department of Chemistry, National Central University, Jhongli 32001, Taiwan, R.O.C; \underline{ZHONG WANG}, Shinnecock Building, SUNY Suffolk Eastern Campus, Riverhead NY 11901; JU XIN, ERIC OTRUBA AND TYLER LATSHA, Department of Physics and Engineering Technology, Bloomsburg University, Bloomsburg, PA, 17815.

Optical double resonance measurements of several rovibronic transitions in the b̃^{1}B_{1} ← ã^{1}A_{1} origin band of CH_{2} were reported at the 2008 Symposium.^{a} Although these assignments are unambiguous, the data are sparse and the spectral resolution was limited by the pulsed laser used. We have therefore recorded the complete absorption spectrum between approximately 8200 cm^{−1} and 8400 cm^{−1} at Doppler-limited resolution using an extended cavity diode laser source. This region includes the expected positions of both the K_{a} = 0 ← 1 and K_{a} = 1 ← 0 subbands. Although the calculated intensities for these transtions are small,^{b} the observed spectra have good signal-to-noise ratio and many additional transitions are seen. We have assigned all the expected transitions involving Ka^{′} = 0, and part of the subband with Ka^{′} = 1. However definitive assignment of transitions involving J′ = 4 and higher in the second subband have so far eluded us, and additional optical double resonance measurements are needed to decide between several possible assignments. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences. Support from the Department of Energy Faculty and Student Teams Program for Bloomsburg University is gratefully acknowledged.

aZ. Wang et al., paper WG03, 63rd OSU International Symposium on Molecular Spectroscopy.

bJ.-P. Gu et al., J. Molec. Struct. 517-8 247 (2000).